OpenStudy (anonymous):

What is the vertex of the graph of the function y=2(x-4)^2+6? Please help fast!

OpenStudy (saifoo.khan):

y=2(x-4)^2+6 Compare it with: y = a(x-h)^2 + k Where the vertex is: (h, k)

OpenStudy (saifoo.khan):

Hence, in this case the vertex will be: (4, 6)

OpenStudy (anonymous):

not -4,6?

OpenStudy (anonymous):

h is the opposite but k stays the same

OpenStudy (anonymous):

oh ok thanks

OpenStudy (saifoo.khan):

Sorry PC crashed. And yes, evy is correct